# A quantitative analysis of Galilei’s observations of Jupiter satellites from the Sidereus Nuncius

### v1

@@ 3 Digitization of the sketches
Tables 1 and 2 summarize the measurements; the first table refers to the observations performed in January while the second to those of February and March. Data have been obtained from the analysis of digitized images using cursors that were manually centred on the satellites. For each image the size of the planet disk was also recorded to normalize the elongations in terms of this diameter. The diameter has always been found to be constant between the sketches with few exceptions (60, 61, 62, 64) where also the nearby star is drawn. The first column is a sequential identifier that correspond to the one of Fig. 1. The following three columns give the time and hour of the observation as reported by Galilei (in italic hours). The 5th column reports the number of visible satellites. The last columns show the position of the satellites ($x_{i}$, $y_{i}$) where $x$ runs horizontally from left to right and $y$ vertically from bottom to top. In the sketches the East is left and West right. This is the most natural choice but is also motivated by the fact that Galilei’s telescope was not inverting the image being composed of a semi-convex or convex objective and bi-concave or plano-concave eyepiece.
@@
A graphical summary of the digitized positions of the satellites, normalized to the diameter of Jupiter as in the previous tables is shown in Fig. 3. The time flows from top to bottom. Comparing this plot with Figs. 1 and 2 ensures the absence of mistakes in the digitization process.
Figure 3: Reconstructed positions of Jupiter satellites from digitization of the sketches in units of the planet’s equatorial diameter. Times increases from top to bottom.
@@ 3.1 Conversion of the observation times
Galilei expressed the observation times in terms of hours since sunset following italic hours that was used at those times. The first hour, for example, refers to one hour after sunset. To convert these data into modern hours (CET) we considered the time of sunset in Padua in the beginning of the year (see Fig. 86, Appendix) and sum the hour reported by Galilei to this value assuming that each hour has the same length. The time of sunset was approximate linearly as: $h_{sunset}=16.58+0.02353~{}d$ where hours are expressed in decimal fractions and $d$ is the ordinal day since the beginning of the year. The result of the conversion is shown in Tab. 3 and 4 where all the observations are associated to a time expressed in CET. It should be noted that 11 times Galilei does not mention the time of the observations (marked as “ïn the column 4 Tab.3 and 4). In this case for the subsequent analysis we assumed that the time was the first hour, a frequent choice. This introduces for these observations an error up to potentially 5-6 hours [^5] The latest observations are before 1 AM. Note that the setting time of Jupiter was about 5.30 and 2.00 at the beginning and end of the considered period. which can be relevant for satellites passing close to Jupiter.
id month day after sunset (h) sunset modern time (h:min)
1 1 7 1:00 16.74 17:44
2 1 8 1:00 16.77 17:46
3 1 10 1:00 16.81 17:48
4 1 11 1:00 16.84 17:50
5 1 12 1:00 16.86 17:51
6 1 13 1:00 16.89 17:53
7 1 15 3:00 16.93 19:55
8 1 15 7:00 16.93 23:55
9 1 16 1:00 16.96 17:57
10 1 17 0:30 16.98 17:28
11 1 17 5:00 16.98 21:58
12 1 18 0:20 17.00 17:20
13 1 19 2:00 17.03 19:01
14 1 19 5:00 17.03 22:01
15 1 20 1:15 17.05 18:17
16 1 20 6:00 17.05 23:02
17 1 20 7:00 17.05 24:02
18 1 21 0:30 17.07 17:34
19 1 22 2:00 17.10 19:05
20 1 22 6:00 17.10 23:05
21 1 23 0:40 17.12 17:47
22 1 23 5:00 17.12 22:07
23 1 24 1:00 17.14 18:08
24 1 24 6:00 17.14 23:08
25 1 25 1:40 17.17 18:50
26 1 26 0:40 17.19 17:51
27 1 26 5:00 17.19 22:11
28 1 27 1:00 17.21 18:12
29 1 30 1:00 17.29 18:17
30 1 31 2:00 17.31 19:18
31 1 31 4:00 17.31 21:18
Table 3: Conversion from Italic to modern hours (January 1610).
id month day after sunset (h) sunset modern time (h:min)
32 2 1 2:00 17.33 19:19
33 2 2 1:00 17.36 18:21
34 2 2 7:00 17.36 24:21
35 2 3 7:00 17.38 24:22
36 2 4 2:00 17.40 19:24
37 2 4 7:00 17.40 24:24
38 2 6 1:00 17.45 18:26
39 2 7 1:00 17.47 18:28
40 2 8 1:00 17.50 18:29
41 2 9 0:30 17.52 18:01
42 2 10 1:30 17.54 19:02
43 2 11 1:00 17.57 18:34
44 2 11 3:00 17.57 20:34
45 2 11 5:30 17.57 23:04
46 2 12 0:40 17.59 18:15
47 2 13 0:30 17.61 18:06
48 2 15 1:00 17.66 18:39
49 2 15 5:00 17.66 22:39
50 2 15 6:00 17.66 23:39
51 2 16 6:00 17.68 23:41
52 2 17 1:00 17.71 18:42
53 2 18 1:00 17.73 18:43
54 2 18 6:00 17.73 23:43
55 2 19 0:40 17.76 18:25
56 2 21 1:30 17.80 19:18
57 2 25 1:30 17.90 19:23
58 2 26 0:30 17.92 18:25
59 2 26 5:00 17.92 22:55
60 2 26 5:00 17.92 22:55
61 2 27 1:04 17.94 19:00
62 2 28 1:00 17.97 18:58
63 2 28 5:00 17.97 22:58
64 3 1 0:40 17.99 18:39
Table 4: Conversion from Italic to modern hours (February, March 1610).
3.2 Satellite-untagged analysis
Before trying to associate the observations to individual satellites, we perform a global analysis putting all the measurements together. The disentangling of the contributions of single satellites was not obvious for Galilei as the data were taken sparsely and often satellites were too close and unresolved. The oscillatory pattern of Callisto is nevertheless quite easily separable (see Fig.  3).
We first check if the distribution is consistent with the sum of four harmonic oscillators corresponding to four distinct circular orbits and then consider how the four frequencies might emerge using the Lomb-Scargle algorithm ([4, 5]).
The distribution of the horizontal displacements ($x_{i}$) is shown in Fig. 4 for the data (bullets) and the expectation in the hypothesis that the motion is harmonic. As expected the probability of observing the satellite when it is at the maximal elongation is higher as its projected velocity is smaller. Galileo did not arrive to notice that the satellites were moving faster when they were closer to Jupiter which is something that might have been at hand with his data.
Figure 4: Distribution of the elongation of the satellites for data (bullets) and a simulation assuming harmonic motions with (green) and without (magenta) the inefficiency effect in the data related to the proximity to Jupiter’s disk.
The expectation in the solid magenta line assumes that satellites are only invisible if they are behind or in front of the disk of Jupiter. The expectation is solid green line shows instead the distribution assuming that the inefficient angular region close to Jupiter corresponds to 2.5 times the real dimension of the disk. It is visible how this second hypothesis is quite consistent with the data.
The superposition of the distribution (that are each normalized to each other) was done by eye by scaling the $x$-axis to match the most extreme observations of Callisto. The dimensions of the orbits are based on modern data. It should be noted that this is very approximate as it assumes that data were taken approximately uniformly in time while, as we saw, Galilei was often taking several, near in time, measurements per night. The large statistical error allows to neglect this and we can say that the observation are semi-qualitatively consistent with what we expect. We will perform fully quantitative considerations anyway in the following.
To isolate the outermost satellite, Callisto, one could only select data being more than four diameters away from Jupiter (see Fig. 4) where only Callisto can reach. After fitting these data with a sinusoid, the points compatible with the fitted Callisto trajectory could be then removed in the innermost part. There one would have been left with three contributions and iterate. Yet looking at the data it is clear that this was not at all straightforward.
@@ 3.3 Lomb-Scargle periodograms
We now try to use the modern technique proposed by Lomb-Scargle [4, 5] to single out the emergence of the expected four distinct mono-chromatic frequencies in the data.
We show in Fig. 5 the LS periodogram of the full dataset.
@@
3.4 Association of the satellites using a sky simulator
This association allows inspecting the expected pattern of satellites using the Stellarium sky simulator.
An example of such an exercise is shown in Fig. 6 for the observation number 7, performed on Jan. 15 1610 at the third hour (Galilei’s time) or 19.55 CET. The correspondence is remarkable and this happens for the majority of the cases as it will be shown later (see Appendix A). It is worth noticing also the level of attention of the recordings where the slight displacement of Callisto from the ecliptic plane is reported.
It is interesting to notice that, according to the simulator, the satellites’ magnitudes were, in decreasing brightness: Ganymede (4.75), Io (5.11), Europa (5.43) and Callisto (5.69). Galilei reports instead in decreasing brightness Callisto, Ganymede, Europa, Io. It is possible that the closeness to Jupiter might have played a role in biasing the perception of the brightness [^6] It should also be noted that the accuracy of the predictions of Stellarium have not been cross-checked with other simulators..
Another evident feature is the fact that the size of Jupiter’s disk is generally overestimated by Galilei. This is probably due to the limited resolution of its telescope. This interpretation is corroborated by the fact the often satellites were missed when too close to the planetary disk, as it will be detailed later.

### v2

@@ 2.1 Dataset from the digitization of the sketches
Tables 2 and 3 summarize the measurements; the first table refers to the observations performed in January while the second to those of February and March. Data have been obtained from the analysis of digitized images using cursors that were manually centered on the satellites. For each image the size of the planet disk was also recorded to normalize the elongations in terms of this diameter. The diameter has always been found to be constant between the sketches with few exceptions (60, 61, 62, 64) where also the nearby star is drawn. The first column is a sequential identifier that correspond to the one of Fig. 2. The following three columns give the time and hour of the observation as reported by Galilei (in italic hours). The 5th column reports the number of visible satellites. The last columns show the position of the satellites ($x_{i}$, $y_{i}$) where $x$ runs horizontally from left to right and $y$ vertically from bottom to top. The coordinates start here at the center of the disk and are normalized to the diameter of the disk in the drawings. In the sketches the East is left and West right. This is the most natural choice but is also motivated by the fact that Galilei’s telescope was not inverting the image being composed of a semi-convex or convex objective and bi-concave or plano-concave eyepiece.
@@ 2.2 Conversion of the observation times
Galilei expressed the observation times in terms of hours since sunset following italic hours that was used at those times. The first hour, for example, refers to one hour after sunset. To convert these data into modern hours (CET) we considered the time of sunset in Padua in the beginning of the year and sum the hour reported by Galilei to this value assuming that each hour has the same length. The time of sunset was approximated linearly as: $h_{sunset}=16.58+0.02353~d$ where hours are expressed in decimal fractions and $d$ is the ordinal day since the beginning of the year. The result of the conversion is shown in Tab. 11 in Appendix where all the observations are associated to a time expressed in CET. The first hour corresponded to about 17:45 CET on Jan. 7 and at about 19:00 at the end of February. The latest observations (7th hour) corresponded instead to about 23:45, at the beginning of January, and 1:00 AM, at the beginning of March. Note that the setting time of Jupiter was about 5.30 and 2.00 at the beginning and end of the considered period.
It should be noted that for 11 observations Galilei does not mention the time. They are marked with a back-slash in the fourth column of Tab. 2 and 3. These are observations 2, 3, 4, 6, 23, 33, 38, 39, 48, 53 and 64. In this case for the subsequent analysis we assumed that the observation time was the first hour, a frequent choice. This introduces for these observations an error up to potentially up to 6 hours in the worst case. This indetermination on the time introduces a significant angular bias especially for satellites close to Jupiter when their angular velocities are of about 20, 15, 12, 8′′/h for Io, Europa Ganymede and Callisto, respectively. Given the angular diameter of Jupiter (45-38′′ in those days) it means that a satellite (especially Io) can easily be visible on different sides of the planet in a six hour time span. In the first column of Tab. LABEL:tab:stellarium in Appendix we have clearly identified these cases with a $\star$. Looking at these eleven configurations it does not seem that this approximation poses specific problems in the interpretation but in some cases it might be argued that some satellites might have been missed because at the real observation hour the satellite might have been closer to Jupiter than at the assumed time (i.e. for obs. 23 Io was not seen even if it seems far away enough for having been seen; in obs. 38 Callisto is not separated from Jupiter even if it seems far enough; similar consideration for Io in obs. 39, Callisto in obs. 48). These cases where hence marked separately in the analysis of the efficiency of the detection near the disk that will be discussed later.
2.3 Dataset from the angular readings of Galileo
In the Sidereus, starting from Jan. 12, Galilei begins to report the positions of the satellites in numbers by providing their angular displacements besides the sketches. He states: It was therefore established by me, and concluded beyond all doubt, that there are in the heavens three wandering stars revolving around Jupiter, similar to Venus and Mercury around the Sun; which was finally observed more clearly than in broad daylight through many subsequent observations: and not only three, but four were found to be the wandering stars performing their revolutions around Jupiter … And I also measured the intervals between them with the method explained earlier. He also says the he noted down the time as he realized how fast were the changes (but still they are missing in 7 cases). Another step in awareness of the impact of the observations happened on Jan. 15, late night, when he started to take notes directly in Latin instead of Italian in view of a publication.
The readings are given in primes (pr.) and seconds of arc (sec). The smallest granularity used is 10′′ but typically distances are expressed in round primes or 1/2 or 1/3 or a prime. Vertical displacements are reported but not quantified. The distant satellites are often characterized by their distance from Jupiter while the other by their relative separations. For a modern reader it is interesting to notice how it was more natural in those days to use an accurate and descriptive account in words with a not fixed pattern rather than choosing to provide a table. Thanks to the precise wording it is anyway easy to do the trivial math and come out with a table where the distances are expressed symmetrically with respect to Jupiter with eastern satellites having negative elongations. The collection of all the measurements is in Tab. 4 (January) and Tab. 5 (February-March).
idid y m d h min $n_{sat}$ $x_{1}$ () $x_{2}$ () $x_{3}$ () $x_{4}$ ()
1 1610 1 7 1 0 0 / / / /
2 1610 1 8 -1 -1 0 / / / /
3 1610 1 10 -1 -1 0 / / / /
4 1610 1 11 -1 -1 0 / / / /
5 1610 1 12 3 0 0 / / / /
5.1 1610 1 12 1 0 2 -2.00 2.00 / /
6 1610 1 13 -1 -1 4 -2.00 1.00 2.00 3.00
7 1610 1 15 3 0 4 2.00 4.00 6.00 11.00
8 1610 1 15 7 0 3 3.00 4.50 9.00 /
9 1610 1 16 1 0 3 -0.67 0.67 8.00 /
10 1610 1 17 0 30 2 -3.00 11.00 / /
11 1610 1 17 5 0 3 -3.00 -2.67 11.00 /
12 1610 1 18 0 20 2 -8.00 10.00 / /
13 1610 1 19 2 0 3 -6.00 5.00 9.00 /
14 1610 1 19 5 0 4 -6.00 -3.00 5.00 9.00
15 1610 1 20 1 15 3 -1.00 1.00 2.00 /
16 1610 1 20 6 0 3 -2.00 0.67 1.00 /
17 1610 1 20 7 0 3 -2.00 0.33 0.83 1.00
18 1610 1 21 0 30 4 -2.50 -1.67 -0.83 4.00
19 1610 1 22 2 0 4 -5.00 1.00 1.67 7.00
20 1610 1 22 6 0 4 -5.00 1.33 2.67 4.00
21 1610 1 23 0 40 3 -7.00 -1.33 3.33 /
22 1610 1 23 5 0 0 / / / /
23 1610 1 24 -1 -1 3 -11.50 -2.50 -2.00 /
24 1610 1 24 6 0 2 -8.00 -3.00 / /
25 1610 1 25 1 40 2 -11.00 -6.00 / /
26 1610 1 26 0 40 3 -11.33 -5.33 5.00 /
27 1610 1 26 5 0 4 -11.33 -5.33 -0.50 5.00
28 1610 1 27 1 0 1 -7.00 / / /
29 1610 1 30 1 0 3 -2.50 3.00 4.00 /
30 1610 1 31 2 0 3 -2.83 -2.33 10.00 /
31 1610 1 31 4 0 3 -2.67 -2.33 10.00 /
32 1610 2 1 2 0 3 -6.00 -0.33 8.00 /
Table 4: Angular displacements data as reported in the text by Galilei starting from Jan. 12 1610. January data.
idid y m d h min $n_{sat}$ $x_{1}$ () $x_{2}$ () $x_{3}$ () $x_{4}$ ()
33 1610 2 2 -1 -1 3 -6.00 4.00 12.00 /
34 1610 2 2 7 0 4 -5.67 -1.67 6.00 14.00
35 1610 2 3 7 0 3 -1.50 2.00 12.00 /
36 1610 2 4 2 0 4 -3.67 -0.67 4.00 10.00
37 1610 2 4 7 0 4 -2.50 -2.00 4.00 7.00
38 1610 2 6 -1 -1 2 -2.00 5.00 / /
39 1610 2 7 -1 -1 2 -2.00 -1.00 / /
40 1610 2 8 1 0 4 -5.67 -5.33 -1.50 -1.33
40.1 1610 2 8 3 0 4 -6.33 -6.00 -1.50 -0.17
40.2 1610 2 8 4 0 3 -6.33 -6.00 -1.50 /
41 1610 2 9 0 30 3 -11.00 -7.00 4.00 /
42 1610 2 10 1 30 2 -10.00 -0.33 / /
42.1 1610 2 10 4 0 1 -12.00 / / /
43 1610 2 11 1 0 3 -12.00 -4.00 4.00 /
44 1610 2 11 3 0 4 -12.00 -4.00 -0.50 4.00
45 1610 2 11 5 30 4 -12.00 -4.00 -2.00 4.00
46 1610 2 12 0 40 4 -10.00 -0.67 1.00 8.00
46.1 1610 2 12 4 0 3 -10.00 1.00 8.00 /
47 1610 2 13 0 30 4 -6.00 -2.00 3.50 4.00
48 1610 2 15 -1 -1 3 -3.17 -1.17 -0.83 /
49 1610 2 15 5 0 2 -4.00 -0.50 / /
50 1610 2 15 6 0 3 -4.00 -0.50 2.00 /
51 1610 2 16 6 0 3 -7.00 5.00 8.00 /
52 1610 2 17 1 0 2 -3.00 10.00 / /
52.1 1610 2 17 6 0 2 -0.83 12.00 / /
53 1610 2 18 1 0 3 -3.00 2.00 10.00 /
53.1 1610 2 18 2 0 3 -3.00 3.00 10.00 /
54 1610 2 18 6 0 4 -4.83 -1.83 3.00 10.00
55 1610 2 19 0 40 2 7.00 13.00 / /
56 1610 2 21 1 30 3 -2.00 3.00 10.00 /
57 1610 2 25 1 30 3 -8.00 -4.00 2.00 /
58 1610 2 26 0 30 2 -10.00 6.00 / /
59 1610 2 26 5 1 3 -11.00 1.00 6.00 /
60 1610 2 26 5 2 0 / / / /
61 1610 2 27 1 4 4 -10.00 -0.50 2.50 3.50
61.1 1610 2 27 5 0 4 -10.00 -1.00 2.50 3.50
62 1610 2 28 1 0 2 -9.00 2.00 / /
63 1610 2 28 5 0 3 -9.00 -2.00 2.00 /
64 1610 3 1 0 40 4 -7.33 -3.33 -3.00 -2.00
64.1 1610 3 2 0 30 3 -7.00 -6.50 2.00 /
Table 5: Angular displacements data as reported in the text by Galilei starting from Jan 12 1610. February-March data.
The first column of Tabs. 4 and 5 are the measurement identifiers already introduced in Tabs. 2 and 3. In this case we have added 9 measurements that are reported in the text but do not have a corresponding sketch. We have denoted them as decimals (5.1, 40.1, 40.2, 42.1, 46.1, 52.1, 53.1, 61.1, 64.1). It is important to notice that the measurements that we are considering are in total 72 (64 with sketches+ 9 with only quoted measurements in the text).
2.4 Comparison of the two datasets
A graphical summary of the positions of the satellites is shown in Fig. 5. The time flows from top to bottom. Comparing this plot with Figs. 2 and 3 ensures the absence of mistakes in the digitization process.
The left plot refers to the dataset with angular information (elongations in primes) while the one on the right comes from the digitization of the sketches and here the measurement are normalized to the diameter of Jupiter as in the drawings of the Sidereus.
Figure 5: Left: reconstructed positions of Jupiter satellites from the angular data of Galilei (in primes). Right: reconstructed positions from the digitization of the sketches in units of the planet’s equatorial diameters in the sketches themselves. Times increases from top to bottom.
In Fig. 6 we show a scatter plot of the elongations with dataset-1 (digitization) and dataset-2 (angular measurements by Galilei). Initially we had interpreted the angular measurements as referred to the center of Jupiter while they should be clearly referred to the limb of Jupiter. In the first, wrong assumption, a clear offset between the positive and negative arms is evident. We have hence corrected the data by adding a half equatorial diameter of Jupiter with a scaling factor that was determined from the data:
$\theta_{i}^{corr}=\theta_{i}+\rm{sign}(\theta_{i})\epsilon\frac{\Delta_{avg}}{2}$ (1)
For the diameter of Jupiter ($\Delta_{avg}$) we have taken the average of the two extreme angular diameters of 45 and 38′′. We have hence varied $\epsilon$ until the minimum $\chi^{2}$ was reached when performing a linear fit of all the points. The best match is achieved for $\epsilon=1.76$ (Fig. 7). This seems to suggest that the diameter of Jupiter was perceived with an expansion factor of this size i.e. of about 79-67′′. The correlation plot using the corrected angular elongations is shown in Fig. 6. The horizontal blue lines indicate the real diameter of Jupiter (38-45′′). The vertical gray lines the disk of Jupiter in the sketches that by definition has a width of one Jupiter diameter. From this figure it become clear how close the satellite recording fall from the disk of Jupiter. The linear fit yields the following relation between the datasets:
$\theta_{i}^{corr}=(1.825\pm 0.06)^{\prime}\times x_{i}+(0.137\pm 0.060)^{\prime}$ (2)
where we denoted with $x_{i}$ the elongations normalized to the diameter of Jupiter as reported in the drawings (dataset-1) and with $\theta_{i}^{corr}$ the angular recordings of Galilei corrected from the limb of Jupiter disk after the tuning done from the data.
Figure 6: Scatter plot of the elongations with dataset-1 (digitization) and dataset-2 (angular measurements by Galilei) after the correction of Eq. 1.
Figure 7: Variation of the $\chi^{2}$ of a linear fit to the data of Fig. 6 as a function of $\epsilon$ in Eq.1, here denoted as “expand”.
In Fig. 8 the dataset-1 and dataset-2 have been superimposed with a proper scaling of dataset-1 given by the the linear fit parameters shown in Fig. 6 and Eq. 2. The entries where only the dataset-1 data appear are those for which Galileo on report the angular data but no additional sketches.
Figure 8: Dataset-1 (black markers) and Dataset-2 (red markers) superimposed after a proper scaling. The entries where only the dataset-1 data appear are those for which Galileo on report the angular data but no additional sketches. The dataset from digitization was expressed in terms of angular displacements using Eq. 2.
2.5 Overall distribution and four harmonic oscillators model
We now check if the distribution of the elongations is consistent with the sum of four harmonic oscillators corresponding to four distinct almost circular orbits. The distribution of the horizontal displacements ($x_{i}$ and $\theta_{i}^{corr}$) is shown in Fig. 9 (left and right plot respectively) for the data (bullets) and the expectation in the hypothesis that the motion is harmonic and neglecting the inclination of the orbits. Here we also implicitly assume that the sampling in time is uniform while in reality we know that observations appeared in clusters during the same night. The expected distribution in the case of the $\theta_{i}^{corr}$ is more smeared than for the $x_{i}$ because we are taking into account in the simulation the variation of the angular diameter of Jupiter from 45 to 38′′. In the diameter normalized dataset naturally this effect is canceled out in the ratio. As expected the probability of observing the satellite when it is at the maximal elongation is higher as its projected velocity is smaller.
Figure 9: Distribution of the elongation of the satellites for data (bullets) and a simulation assuming harmonic motions with (green) and without (magenta) the inefficiency effect in the data related to the proximity to Jupiter’s disk. Left plot refers to dataset-1, right plot to dataset-2.
The expectation in the solid magenta line assumes that satellites are only invisible if they are behind or in front of the disk of Jupiter. The expectation is solid green line shows instead the distribution assuming that the inefficient angular region close to Jupiter corresponds to 2.5 and 1.76 times the real dimension of the disk. These hypotheses are quite consistent with the data. The factor 1.76 for the angular data derives from the analysis presented before. Further discussion of this inefficiency near the disk is presented in Sec. 5.
For dataset-1 the superposition of the distribution (that are each normalized to each other) was done by eye by scaling the $x$-axis to match the most extreme observations of Callisto. For dataset-2 the data had to be scaled down by a factor 0.7 to match with modern expectations on the satellites’ angular elongations during that observational period. This systematic effect will be discussed in more detail in the context of the sinusoidal fits later in Sec. 4.2.
Despite the mentioned approximations, the observation are semi-qualitatively consistent with what we expect, also considering the large statistical uncertainties. We will perform fully quantitative considerations on the dataset where the information of individual satellites is extracted (Sec. 4.2).
3 Satellite-untagged analysis
Before trying to associate the observations to individual satellites, we perform a global analysis considering all the measurements together. The disentangling of the contributions of single satellites was a rather hard task as the data were taken sparsely and often satellites were too close and unresolved. This was the subject of a significant effort by Galilei until 1619[24] and several other astronomers (i.e. mainly Pereisc, Harriot, Mayr but also Kepler, Magini, Sizzi, Agucchi) as reported in [6, 16]. It was soon understood how important this could be as a tool for determining the longitude on Earth. The “atlantic labour” of Galilei is analysed in a publication by Stillman Drake[11]. It is interesting to notice anyway that the oscillatory pattern of Callisto is already quite easily separable when the elongation is sufficiently large (see Fig.  10). Apart from this sorting put a pattern from the data is rather hard. To isolate the outermost satellite, Callisto, one could only select data being more than four diameters away from Jupiter (see Fig. 9) where only Callisto can reach. After fitting these data with a sinusoid, the points compatible with the fitted Callisto trajectory could be then removed in the innermost part. There one would have been left with three contributions and iterate. Yet looking at the data it is clear that this was not at all straightforward.
We now consider how the four frequencies might emerge using the Lomb-Scargle algorithm which is a classical algorithm to single out the emergence of particular frequencies in a discrete and sparsed dataset ([25, 26]).
@@ 3.1 Lomb-Scargle periodograms
We show in Fig. 10 the Lomb-Scargle (LS) periodogram of the full dataset.
@@
4 Association of the satellites using a sky simulator
This association allows inspecting the expected pattern of satellites using the Stellarium [20] online sky simulator.
An example of such an exercise is shown in Fig. 12 for the observation number 7, performed on Jan. 15 1610 at the third hour (Galilei’s time) or 19.55 CET. The correspondence is remarkable and this happens almost for the entire set of observations as it will be shown later (see Appendix A). It is worth noticing also the level of attention of the recordings where the slight displacement of Callisto from the ecliptic plane is reported. The orbits are almost perfectly circular and lying almost exactly at the the Jupiter equatorial plane. The rotation axis of Jupiter forms an angle of 3.1 with respect to the Ecliptic and the inclination of the Jupiter’s orbit forms with it an angle of 1.3. Displacements in the “vertical” direction arise from the combinations of these effects and the satellites are describing very narrow ellipses. Figure 11 shows these trajectories in the considered period as predicted ([1]). The vertical displacement of Callisto near Jupiter amounted to about $\pm$20′′.
Figure 11: Foreseen orbits of the Galileian satellites in the considered period[1].
It is interesting to notice that, according to the simulator, the satellites’ magnitudes were, in decreasing brightness: Ganymede (4.75), Io (5.11), Europa (5.43) and Callisto (5.69). Galilei reports instead in decreasing brightness Callisto, Ganymede, Europa, Io. It is possible that the closeness to Jupiter might have played a role in biasing the perception of the brightness [^3] It should also be noted that the accuracy of the predictions of Stellarium for the magnitudes of the satellites have not been cross-checked with other simulators..
Another evident feature is the fact that the size of Jupiter’s disk is generally overestimated by Galilei. This is probably due to the glare of the planet and resolution effects. This interpretation is corroborated by the fact the often satellites were missed when too close to the planetary disk, as it will be detailed later.